# AutoOPE: Automated Off-Policy Estimator Selection

### v1

@@ Abstract
The Off-Policy Evaluation (OPE) problem consists of evaluating the performance of counterfactual policies with data collected by another one. This problem is of utmost importance for various application domains, e.g., recommendation systems, medical treatments, and many others. To solve the OPE problem, we resort to estimators, which aim to estimate in the most accurate way possible the performance that the counterfactual policies would have had if they were deployed in place of the logging policy. In the literature, several estimators have been developed, all with different characteristics and theoretical guarantees. Therefore, there is no dominant estimator, and each estimator may be the best one for different OPE problems, depending on the characteristics of the dataset at hand. While the selection of the estimator is a crucial choice for an accurate OPE, this problem has been widely overlooked in the literature. We propose an automated data-driven OPE estimator selection method based on machine learning. In particular, the core idea we propose in this paper is to create several synthetic OPE tasks and use a machine learning model trained to predict the best estimator for those synthetic tasks. We empirically show how our method is able to generalize to unseen tasks and make a better estimator selection compared to a baseline method on several real-world datasets, with a computational cost significantly lower than the one of the baseline.
@@ 1 Introduction
In numerous real-world scenarios, decision-making systems are required to interact with dynamic environments, making choices aimed at optimizing specific objectives. These systems are prevalent in domains such as recommender systems [10, 30, 7], information retrieval [22], ad-placement systems [4], or medical treatments [9]. Such decision-making problems can be viewed through the lens of the Contextual Bandit paradigm [43, 19]. For instance, let us consider the recommendation system domain: the recommender (agent) repeatedly observes user profiles (contexts), and aims to recommend items (actions) to users based on the given context to maximize some notion of reward.
A critical challenge arises in evaluating the performance of new policies without deploying them in the real world. This is especially pertinent in scenarios where deploying a suboptimal policy can have significant negative consequences, such as lost revenue in ad placement or user dissatisfaction in recommender systems. Off-Policy Evaluation (OPE) seeks to address this challenge by evaluating the performance of counterfactual policies using data collected from a different, previously deployed policy, often termed the logging policy. The ability to make informed decisions about policy modifications without necessitating real-world deployment is not only cost-effective but also reduces the potential risks associated with untested policies.
@@
However, the effectiveness of OPE largely hinges on the accuracy of the estimator used to predict the performance of the evaluation policies. There is a wide variety of estimators that have been proposed in the literature, each with distinct characteristics and mathematical properties [41]. As an example, a well-known estimator, called Inverse Propensity Scoring (IPS) [29, 34], mitigates the estimation bias by adjusting observed rewards based on the inverse probability assigned by the logging policy to the logged action. In this way, actions that have a low probability under the logging policy receive higher weighting to compensate for their infrequent occurrence in the dataset, and vice versa for actions more likely under the logging policy. Despite IPS’s prominence, numerous other estimators exist, each with unique trade-offs between bias and variance. Given the heterogeneity of real-world datasets and the myriad of scenarios in which OPE is applied, no single estimator is universally optimal. Further details on related work are reported in Appendix A. An estimator may be the best under certain circumstances but perform poorly under others. This variety of estimators, while offering flexibility, also introduces the challenge of Estimator Selection, which is a crucial problem in OPE and has been mostly overlooked in the existing literature. Basically, the Estimator Selection problem consists of the following: we are given an Off-Policy Evaluation task and we have to select the best estimator for the task we have at hand.
@@
In this paper, we address the Estimator Selection problem using an automated data-driven approach. Our main idea consists of exploiting synthetic data. We generate several different synthetic OPE tasks by varying the characteristics of each one. Various types of estimators are applied to these tasks, and their estimates are compared against the ground-truth value of each evaluation policy, which is known due to the synthetic nature of the data. These synthetic OPE tasks are used to train a machine learning model designed to predict estimator’s Mean Squared Errors for the given OPE task. This model can select the estimator with the lowest predicted error, effectively addressing the OPE Estimator Selection challenge. We call our Estimator Selection method Automated Off-Policy Estimator Selection (AutoOPE). The problem of Estimator Selection has been widely overlooked in the literature, and, to the best of our knowledge, only a very recent paper [40] proposed a method for Estimator Selection, named Policy-Adaptive Estimator Selection via Importance Fitting (PAS-IF), which is substantially different from our proposal.
@@
  • To the best of our knowledge, this is the first work to address the estimator selection problem in OPE using a data-driven approach leveraging only synthetic data, and it is one of the few techniques proposed for Estimator Selection in OPE in general, despite the importance of the problem.
  • We conduct extensive experiments on several datasets that empirically demonstrate the effectiveness of our proposed AutoOPE method in selecting high-performing estimators tailored to the given OPE task, performing better than the baseline with a much reduced computational cost. This represents an important contribution towards automated and adaptive Off-Policy Evaluation.
@@ 2.1 Contextual Bandit Problem
The Contextual Bandit (CB) problem is an instance of sequential decision-making problems. At a given time instant $t$, an agent observes a context $x_{t}\sim p(\cdot)$ from a space $\mathcal{X}$, and chooses an action $a_{t}$ from a set $\mathcal{A}$, according to a policy $\pi_{b}$, that is a statistical distribution of the actions conditioned on the context ($a_{t}\sim\pi_{b}(\cdot|x_{t})$), called logging policy. After choosing an action $a_{t}$ the agent receives a reward $r_{t}$ from the environment, sampled from $r_{t}\sim p(\cdot|x_{t},a_{t})$. This means that the reward observed is conditioned on both context and action.
A common quantity used to compare the quality of different policies is the policy value, that is defined as the expected value of the reward collected under a certain policy $\pi$: $V(\pi)\coloneqq\mathbb{E}_{p(r_{t}|x_{t},a_{t})\pi(a_{t}|x_{t})p(x_{t})}\left[r_{t}\right]$.
@@ 2.2 Off-Policy Evaluation
A common metric used to assess the performance of an estimator is the Mean Squared Error (MSE), defined as $MSE(\hat{V}(\pi_{e}))=\mathbb{E}\left[(\hat{V}(\pi_{e})-V(\pi_{e}))^{2}\right]$.
@@ 2.3 Estimator Selection for OPE
The effectiveness of OPE crucially depends on how accurate is the estimator in estimating the policy value of the evaluation policy, usually in terms of how low is its $MSE(\hat{V}(\pi_{e}))$. For this reason, and since many OPE estimators have been proposed in the literature, selecting the best estimator for a given OPE task, i.e., Estimator Selection, is particularly important. Given a finite set $\mathcal{V}$ of different OPE estimators, the Estimator Selection problem consists in finding the best estimator among them.
Since in OPE we do not have any data collected under the evaluation policy, we cannot compute the ground-truth evaluation policy value $V(\pi_{e})$ and therefore the ground-truth MSE of the estimator. A common approach is to approximate it with an estimate, i.e., $\widehat{MSE}(\hat{V}(\pi_{e}))$, and to select the estimator that has the lowest approximate MSE.
@@ 2.3.1 Policy-Adaptive Estimator Selection via Importance Fitting (PAS-IF)
To the best of our knowledge, only Policy-Adaptive Estimator Selection via Importance Fitting (PAS-IF) [40] has tackled the problem of Estimator Selection in the literature. To accurately estimate the MSE of an OPE estimator, the key idea of PAS-IF is to determine a sampling rule $\rho_{\theta}:\mathcal{X}\times\mathcal{A}\rightarrow(0,1)$, that governs a subsampling procedure to divide the logging data $\mathcal{D}_{b}$ into two pseudo datasets: a pseudo logging dataset $\tilde{\mathcal{D}}_{b}$ and a pseudo evaluation dataset $\tilde{\mathcal{D}}_{e}$. $\rho_{\theta}(x_{t},a_{t})$ represents the probability of assigning the logging sample $(x_{t},a_{t})$ to $\tilde{\mathcal{D}}_{e}$. In this context, the two pseudo datasets can be viewed as if they were collected by two pseudo policies, $\tilde{\pi}_{b}$ and $\tilde{\pi}_{e}$, induced by $\rho_{\theta}$. The $\rho_{\theta}(x_{t},a_{t})$ is the output of a fully connected neural network, optimized by minimizing the average squared error between the importance ratio $w(x_{t},a_{t})=\pi_{e}(a_{t}|x_{t})/\pi_{b}(a_{t}|x_{t})$ induced by true policies, and $\tilde{w}(x_{t},a_{t})=\tilde{\pi}_{e}(a_{t}|x_{t})/\tilde{\pi}_{b}(a_{t}|x_{t})$ induced by the pseudo policies: this learning phase is called importance fitting. The pseudo evaluation dataset $\tilde{\mathcal{D}}_{e}$ is deemed to be collected by the evaluation policy, and in this way it can be used to estimate the evaluation policy value by averaging the rewards in it (we denote this estimate as $\hat{V}_{on}(\pi_{e};\mathcal{\tilde{D}}_{e})$). The pseudo logging dataset $\tilde{\mathcal{D}}_{b}$ is used to compute the candidate OPE estimator prediction ($\hat{V}(\pi_{e};\mathcal{\tilde{D}}_{b})$). After that, each estimator MSE can be approximated in this way:
$\widehat{MSE}_{\mbox{PAS-IF}}(\hat{V}(\pi_{e}))\coloneqq(\hat{V}_{on}(\pi_{e};\mathcal{\tilde{D}}_{e})-\hat{V}(\pi_{e};\mathcal{\tilde{D}}_{b}))^{2}\ .$
@@ 3.1 Theoretical Motivation
AutoOPE is theoretically justified by many findings that show how the performance of an OPE estimator is highly dependent on the characteristics of the OPE task at hand [42, 32]. For instance, [24, 25, 26] theoretically analyze Inverse Propensity Scoring (IPS) [29, 34], one of the most used OPE estimators and the basis for many others. Metelli et al. [24] show that the IPS error is strictly related to the (exponentiated) Rényi divergence between the logging and evaluation policies [^2] The Rényi divergence is a divergence between two probability distributions, and it is defined as $d_{r}(\pi_{e}||\pi_{b})=\log\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{\ a\sim\pi_{b}}\left[(\pi_{e}(a|x)/\pi_{b}(a|x))^{2}\right]$. and provide a finite-sample bound on the absolute estimation error of IPS under standard assumptions:
$\left\lvert\hat{V}_{\text{IPS}}(\pi_{e})-V(\pi_{e})\right\rvert\leq\mathcal{O}\left(\sqrt{\frac{\exp(d_{r}(\pi_{e}||\pi_{b}))}{n_{b}}}\right),$
later also proving that this bound is tight [26]. This theoretical finding, along with many others, shows how the performance of an OPE estimator is strictly related to the characteristics of the OPE task, such as the divergence between the evaluation and the logging policy. In other words, we assume that:
$MSE(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}})=f^{*}(g(\hat{V},\mathcal{T}_{\text{ope}}))\ ,$
@@
which means that there is a function $f^{*}$ that express the relationship between the real error of an estimator on a given OPE task $MSE(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}})$ and the features $g(\hat{V},\mathcal{T}_{\text{ope}})$ of the estimator and the task. For this reason, we propose our zero-shot meta-learning approach [8], which we frame as a supervised learning problem.
@@
Let us assume that we have a large variety of tasks at our disposal $\mathcal{M}\coloneqq\{\mathcal{T}_{\text{ope}}^{(1)},\mathcal{T}_{\text{ope}}^{(2)},\dots\}$, a set of estimators $\mathcal{V}$, and the feature extraction function $g$. Also, let us assume that we have access to the ground-truth MSE. If we have access to these, we could try to find the function that best predicts the MSE for all tasks and estimators, as a standard supervised problem:
$\hat{f}=\arg\min_{f}\sum_{\mathcal{T}_{\text{ope}}\in\mathcal{M}}\sum_{\hat{V}\in\mathcal{V}}\mathcal{L}\left(MSE(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}}),f(g(\hat{V},\mathcal{T}_{\text{ope}}))\right)\ ,$
@@
where $\mathcal{L}$ is a suitable loss. Once we found $\hat{f}$, we can apply it in a zero-shot way whenever we need to perform estimator selection with a new task $\mathcal{T}_{\text{ope}}^{(\text{new})}\notin\mathcal{M}$. Indeed, with $\hat{f}$ we can now estimate the MSE for a given task in this way:
$\widehat{MSE}_{\mbox{AutoOPE}}\left(\hat{V},\mathcal{T}_{\text{ope}}^{(\text{new})}\right)\coloneqq\hat{f}(g(\hat{V},\mathcal{T}_{\text{ope}}^{(\text{new})})).$
@@
$\hat{V}^{*}=\arg\min_{\hat{V}\in\mathcal{V}}\widehat{MSE}_{\mbox{AutoOPE}}\left(\hat{V},\mathcal{T}_{\text{ope}}^{(\text{new})}\right).$

### v2

# Automated Off-Policy Estimator Selection via Supervised Learning

@@ Abstract
The Off-Policy Evaluation (OPE) problem consists of evaluating the performance of counterfactual policies with data collected by another one. To solve the OPE problem, we resort to estimators, which aim to estimate in the most accurate way possible the performance that the counterfactual policies would have had if they were deployed in place of the logging policy. In the literature, several estimators have been developed, all with different characteristics and theoretical guarantees. Therefore, there is no dominant estimator and each estimator may be the best for different OPE problems, depending on the characteristics of the dataset at hand. Although the selection of the estimator is a crucial choice for an accurate OPE, this problem has been widely overlooked in the literature. We propose an automated data-driven OPE estimator selection method based on supervised learning. In particular, the core idea we propose in this paper is to create several synthetic OPE tasks and use a machine learning model trained to predict the best estimator for those synthetic tasks. We empirically show how our method is able to perform a better estimator selection compared to a baseline method on several real-world datasets, with a computational cost significantly lower than the one of the baseline.
@@ 1 INTRODUCTION
In numerous real-world scenarios, decision-making systems are required to interact with dynamic environments, making choices aimed at optimizing specific objectives. These systems are prevalent in domains such as recommender systems (Gilotte et al., 2018; Saito and Joachims, 2021; Felicioni et al., 2022), information retrieval (Li et al., 2015), ad-placement systems (Bottou et al., 2013), or medical treatments (Gan et al., 2023). Such decision-making problems can be viewed through the lens of the contextual bandit paradigm (Wang et al., 2005; Langford and Zhang, 2007). For instance, let us consider the recommendation system domain: the recommender (agent) repeatedly observes user profiles (contexts), and aims to recommend items (actions) to users based on the given context to maximize some notion of reward. A critical challenge arises in evaluating the performance of new policies without deploying them in the real world. This is especially pertinent in scenarios where deploying a suboptimal policy can have significant negative consequences. Off-Policy Evaluation (OPE) addresses this challenge by evaluating the performance of counterfactual policies using data collected from a different, previously deployed policy, often termed the logging policy. The ability to make informed decisions about policy modifications without necessitating real-world deployment is not only cost-effective but also reduces the potential risks associated with untested policies.
@@
However, the effectiveness of OPE largely hinges on the accuracy of the estimator used to predict the performance of the evaluation policies. There is a wide variety of estimators that have been proposed in the literature, each with distinct characteristics and mathematical properties (Vlassis et al., 2019). As an example, a well-known estimator, called Inverse Propensity Scoring (IPS) (Precup et al., 2000; Strehl et al., 2010), provides an unbiased estimation, at the cost of high variance. Other estimators also exist, each with unique trade-offs between bias and variance. Given the heterogeneity of real-world datasets and the myriad of scenarios in which OPE is applied, no single estimator is universally optimal. An estimator may be the best under certain circumstances but perform poorly under others. This variety of estimators, while offering flexibility, also introduces the challenge of Estimator Selection, which is a crucial problem in OPE and has been mostly overlooked in the existing literature. Basically, the Estimator Selection problem consists of the following: we are given an Off-Policy Evaluation task and we have to select the best estimator for the task we have at hand.
@@
In this paper, we address the Estimator Selection problem using an automated data-driven approach. Our main idea consists of framing this problem as a supervised learning problem. In particular, we notice that there are several results suggesting that an estimator’s error is related to the OPE task characteristics. Hence, we generate several different synthetic OPE tasks by varying the characteristics of each one, we apply various types of estimators and their estimates are compared against the ground-truth value of each evaluation policy, which is known due to the synthetic nature of the data. These synthetic OPE tasks are used to train a supervised model designed to predict estimator’s Mean Squared Error for the given OPE task. The supervised model will understand the underlying patterns shared among the various OPE tasks and their relation with the estimator error, and then it can be used to select the estimator with the lowest predicted error, effectively addressing the OPE Estimator Selection challenge. We call our Estimator Selection method Automated Off-Policy Estimator Selection (AutoOPE). The problem of Estimator Selection has been widely overlooked in the literature, and, to the best of our knowledge, only a very recent paper (Udagawa et al., 2023) proposed a method for Estimator Selection, named Policy-Adaptive Estimator Selection via Importance Fitting (PAS-IF), which is substantially different from our proposal.
@@
  • To the best of our knowledge, this is one of the few techniques proposed for Estimator Selection, despite the importance of the problem.
  • We conduct extensive experiments on several datasets that empirically demonstrate the effectiveness of our proposed AutoOPE method in selecting high-performing estimators tailored to the given OPE task, performing better than the baseline with a much reduced computational cost.
@@ 2.1 Contextual Bandit Problem
The Contextual Bandit (CB) problem is an instance of sequential decision-making problems. At a given time instant $t$, an agent observes a context $x_{t}\sim p(\cdot)$ from a space $\mathcal{X}$, and chooses an action $a_{t}$ from a set $\mathcal{A}$, according to a policy $\pi_{b}$, that is a statistical distribution of the actions conditioned on the context ($a_{t}\sim\pi_{b}(\cdot|x_{t})$), called logging policy. After choosing an action $a_{t}$ the agent receives a reward $r_{t}$ from the environment, sampled from $r_{t}\sim p(\cdot|x_{t},a_{t})$. This means that the reward observed is conditioned on both context and action. A common quantity used to compare the quality of different policies is the policy value, that is defined as the expected value of the reward collected under a certain policy $\pi$: $V(\pi)\coloneqq\mathbb{E}_{p(r_{t}|x_{t},a_{t})\pi(a_{t}|x_{t})p(x_{t})}\left[r_{t}\right]$.
@@ 2.2 Off-Policy Evaluation
A common metric used to assess the performance of an estimator is the Mean Squared Error (MSE), defined as $\text{MSE}(\hat{V}(\pi_{e}))=\mathbb{E}\left[(\hat{V}(\pi_{e})-V(\pi_{e}))^{2}\right]$.
@@ 2.3 Estimator Selection for OPE
The effectiveness of OPE crucially depends on how accurate is the estimator in estimating the policy value of the evaluation policy, usually in terms of how low is its $\text{MSE}(\hat{V}(\pi_{e}))$. For this reason, and since many OPE estimators have been proposed in the literature, selecting the best estimator for a given OPE task, i.e., Estimator Selection, is particularly important. Given a finite set $\mathcal{V}$ of different OPE estimators, the Estimator Selection problem consists in finding the best estimator among them.
Since in OPE we do not have any data collected under the evaluation policy, we cannot compute the ground-truth evaluation policy value $V(\pi_{e})$ and therefore the ground-truth MSE of the estimator. A common approach is to approximate it with an estimate, i.e., $\widehat{\text{MSE}}(\hat{V}(\pi_{e}))$, and to select the estimator that has the lowest approximate MSE.
@@ 2.3.1 Policy-Adaptive Estimator Selection via Importance Fitting (PAS-IF)
To the best of our knowledge, only Policy-Adaptive Estimator Selection via Importance Fitting (PAS-IF) (Udagawa et al., 2023) has tackled the problem of Estimator Selection in the literature. To accurately estimate the MSE of an OPE estimator, the key idea of PAS-IF is to determine a sampling rule $\rho_{\theta}:\mathcal{X}\times\mathcal{A}\rightarrow(0,1)$, that governs a subsampling procedure to divide the logging data $\mathcal{D}_{b}$ into two pseudo datasets: a pseudo logging dataset $\tilde{\mathcal{D}}_{b}$ and a pseudo evaluation dataset $\tilde{\mathcal{D}}_{e}$. $\rho_{\theta}(x_{t},a_{t})$ represents the probability of assigning the logging sample $(x_{t},a_{t})$ to $\tilde{\mathcal{D}}_{e}$. In this context, the two pseudo datasets can be viewed as if they were collected by two pseudo policies, $\tilde{\pi}_{b}$ and $\tilde{\pi}_{e}$, induced by $\rho_{\theta}$. The $\rho_{\theta}(x_{t},a_{t})$ is the output of a fully connected neural network, optimized by minimizing the average squared error between the importance ratio $w(x_{t},a_{t})=\pi_{e}(a_{t}|x_{t})/\pi_{b}(a_{t}|x_{t})$ induced by true policies, and $\tilde{w}(x_{t},a_{t})=\tilde{\pi}_{e}(a_{t}|x_{t})/\tilde{\pi}_{b}(a_{t}|x_{t})$ induced by the pseudo policies: this learning phase is called importance fitting. The pseudo evaluation dataset $\tilde{\mathcal{D}}_{e}$ is deemed to be collected by the evaluation policy, and in this way it can be used to estimate the evaluation policy value by averaging the rewards in it (we denote this estimate as $\hat{V}_{on}(\pi_{e};\mathcal{\tilde{D}}_{e})$). The pseudo logging dataset $\tilde{\mathcal{D}}_{b}$ is used to compute the candidate OPE estimator prediction ($\hat{V}(\pi_{e};\mathcal{\tilde{D}}_{b})$). After that, each estimator MSE can be approximated in this way:
$\widehat{\text{MSE}}_{\mbox{PAS-IF}}(\hat{V}(\pi_{e}))\coloneqq(\hat{V}_{on}(\pi_{e};\mathcal{\tilde{D}}_{e})-\hat{V}(\pi_{e};\mathcal{\tilde{D}}_{b}))^{2}\ .$
@@ 3.1 Theoretical Motivation
AutoOPE is motivated by many findings that show how the performance of an OPE estimator is highly dependent on the characteristics of the OPE task at hand Voloshin et al. (2021); Saito et al. (2021b). To make one example, Metelli et al. (2018) show that the IPS error is strictly related to the (exponentiated) Rényi divergence (with $\alpha=2$) between the logging and evaluation policies [^2] The Rényi divergence is a family of divergences between two probability distributions with a parameter $\alpha$. For $\alpha=2$, it is defined as $d_{r}(\pi_{e}||\pi_{b})=\log\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{\ a\sim\pi_{b}}\left[(\pi_{e}(a|x)/\pi_{b}(a|x))^{2}\right]$. and provides a finite sample bound on the absolute estimation error of IPS under standard assumptions: $\left\lvert\hat{V}_{\text{IPS}}(\pi_{e})-V(\pi_{e})\right\rvert\leq\mathcal{O}\left(\sqrt{\frac{\exp(d_{r}(\pi_{e}||\pi_{b}))}{n_{b}}}\right).$ This theoretical finding is just one example among many others that shows how the performance of an OPE estimator is related to the characteristics of the OPE task. In other words, we make the following assumption:
Assumption 1.
@@
Consider a fixed set of estimators $\mathcal{V}$. There exists a function $f^{*}\in\mathcal{F}$ that expresses the relationship between the real error of any estimator on a given OPE task $MSE(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}})$ and the features $g(\hat{V},\mathcal{T}_{\text{ope}})$ of the estimator and the task:
$\text{MSE}(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}})=f^{*}(g(\hat{V},\mathcal{T}_{\text{ope}})),\quad\forall\hat{V}\in\mathcal{V}.$
@@
Let us assume that we have a variety of tasks at our disposal $\mathcal{M}\coloneqq\{\mathcal{T}_{\text{ope}}^{(1)},\mathcal{T}_{\text{ope}}^{(2)},\dots\}$, a fixed set of estimators $\mathcal{V}$, and the feature extraction function $g$. With these, we can frame a supervised learning problem and apply empirical risk minimization to find the function that best predicts the MSE for all tasks and estimators:
$\begin{split}\hat{f}=&\arg\min_{f\in\mathcal{F}}\frac{1}{N}\sum_{\begin{subarray}{c}\hat{V}\\ \mathcal{T}_{\text{ope}}\end{subarray}}\mathcal{L}\Bigl{(}\text{MSE}(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}}),f(g(\hat{V},\mathcal{T}_{\text{ope}}))\Bigr{)},\end{split}$
where $\mathcal{L}$ is a suitable loss, $\mathcal{F}$ is the hypothesis class, and $N=|\mathcal{V}||\mathcal{M}|$. Now, we can prove that the empirical risk minimizer $\hat{f}$ is accurate with high probability. To provide an intuition, we consider a finite hypothesis class, but one can extend this finding to infinite hypothesis classes via standard learning theory tools (e.g., VC-dimension, Rademacher complexity) (Shalev-Shwartz and Ben-David, 2014). Also, we consider a subgaussian loss, which is less restrictive than only considering bounded loss functions (Germain et al., 2016).
Assumption 2 (Sub-Gaussian Loss).
The loss function $\mathcal{L}$ is subgaussian with parameter $\sigma^{2}$. Specifically, for all $y,\hat{y}$, the random variable $\mathcal{L}(y,\hat{y})-\mathbb{E}[\mathcal{L}(y,\hat{y})]$ satisfies:
$\mathbb{E}\left[e^{\lambda(\mathcal{L}(y,\hat{y})-\mathbb{E}[\mathcal{L}(y,\hat{y})])}\right]\leq e^{\frac{\lambda^{2}\sigma^{2}}{2}}\quad\forall\lambda\in\mathbb{R}.$
Assumption 3 (Finite hypothesis class).
The hypothesis class $\mathcal{F}$ is finite: $|\mathcal{F}|<\infty$.
Theorem 1.
Under Assumptions 1, 2, and 3, suppose we have $N=|\mathcal{V}||\mathcal{M}|$ independent samples $\{(x_{i},y_{i})\}_{i=1}^{N}$, where each $x_{i}=g(\hat{V}_{i},\mathcal{T}_{i})$ and $y_{i}=MSE(\hat{V}_{i}(\pi_{e});\mathcal{T}_{i})$, drawn uniformly over the estimators $\hat{V}\in\mathcal{V}$ and tasks $\mathcal{T}\in\mathcal{M}$. Let $\hat{f}$ be the empirical risk minimizer over $\mathcal{F}$:
$\hat{f}=\arg\min_{f\in\mathcal{F}}\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}\left(y_{i},f(x_{i})\right).$
Then, for any $\delta>0$, with probability at least $1-\delta$, the expected loss $L(\hat{f})$ satisfies:
$L(\hat{f})\leq\sqrt{\frac{8\sigma^{2}\log\left(\dfrac{2|\mathcal{F}|}{\delta}\right)}{N}},$
where $\sigma^{2}$ is the subgaussian parameter from Assumption 2, and the expected loss $L(f)$ is defined as
$L(f)=\mathbb{E}_{(x,y)}\left[\mathcal{L}\left(y,f(x)\right)\right].$
Proof.
The proof derives from mapping the estimator selection problem to a supervised learning one, and then applying standard learning theory tools, such as Hoeffding’s inequality and the union bound. See Appendix B for all the deferred proofs. ∎
Theorem 1 shows that, with high probability, the empirical risk minimizer $\hat{f}$ has a low error, and the error term that decreases with $N$. Since the number of estimator is fixed, the error term decreases the number of tasks $|\mathcal{M}|$ contained in the dataset.
@@
This theoretical finding is intuitive, and it shows that $\hat{f}$ is trained to understand the patterns of the given tasks and it learns more if provided with more tasks. Once we have found $\hat{f}$, we can apply it in a zero-shot way whenever we need to perform estimator selection with a new task $\mathcal{T}_{\text{ope}}^{(\text{new})}\notin\mathcal{M}$. With $\hat{f}$ we can now estimate the MSE for a given task in this way:
$\widehat{MSE}_{\mbox{AutoOPE}}(\hat{V},\mathcal{T}_{\text{ope}}^{(\text{new})})\coloneqq\hat{f}(g(\hat{V},\mathcal{T}_{\text{ope}}^{(\text{new})})).$
@@
$\hat{V}^{*}=\arg\min_{\hat{V}\in\mathcal{V}}\widehat{MSE}_{\mbox{AutoOPE}}(\hat{V},\mathcal{T}_{\text{ope}}^{(\text{new})}).$